One side of a triangle is half the longest side. The third side is 8 meters less than the longest side. The perimeter is 42 meters. Find all three sides. (Enter your answers from smallest to largest.)
? meters
? meters
? meters

Answer by ewatrrr(24785)


Hi
Let x, x/2, (x-8), represent the sides
Question states***
x + x/2 + (x-8) = 42m
Solving for x
5/2x = 50m
x = 20m x/2 = 10m and (x-8)= 12m
Length of sides are: 10m,12m,20m
CHECKING our Answer***
20m + 10m + 12m = 42m